# Can We Trust AI Agents? An Experimental Study Towards Trustworthy LLM-Based Multi-Agent Systems for AI Ethics

### v1

@@ Abstract
AI-based systems, including Large Language Models (LLMs), impact millions by supporting diverse tasks but face issues like misinformation, bias, and misuse. Ethical AI development is crucial as new technologies and concerns emerge, but objective, practical ethical guidance remains debated. This study examines LLMs in developing ethical AI systems, assessing how trustworthiness-enhancing techniques affect ethical AI output generation. Using the Design Science Research (DSR) method, we identify techniques for LLM trustworthiness: multi-agents, distinct roles, structured communication, and multiple rounds of debate. We design the multi-agent prototype LLM-BMAS, where agents engage in structured discussions on real-world ethical AI issues from the AI Incident Database. The prototype’s performance is evaluated through thematic analysis, hierarchical clustering, ablation studies, and source code execution. Our system generates around 2,000 lines per run, compared to only 80 lines in the ablation study. Discussions reveal terms like bias detection, transparency, accountability, user consent, GDPR compliance, fairness evaluation, and EU AI Act compliance, showing LLM-BMAS’s ability to generate thorough source code and documentation addressing often-overlooked ethical AI issues. However, practical challenges in source code integration and dependency management may limit smooth system adoption by practitioners. This study aims to shed light on enhancing trustworthiness in LLMs to support practitioners in developing ethical AI-based systems.
@@ 1 Introduction
Artificial Intelligence (AI) is emerging as a transformative force, reshaping industries, economies and everyday life. Despite its rapid development and adoption, a number of negative reports on its use highlight the importance of adhering to ethical norms and principles [39]. Several ethical guidelines are available with plenty ethical principles providing high level and abstract guidance for developers and stakeholders on AI ethics [3]. These are important, but there is a lack of practical guidance for developers to operationalise ethics in AI [3]. As new AI-based technologies emerge, ethics in AI will also become increasingly critical, such as the latest breakthrough: Large Language Models (LLMs).
LLMs are becoming ubiquitous and have significant impact on decision-making processes and human interactions [20, 33]. LLMs are advanced AI algorithms capable of generating, interpreting, and predicting text based on vast amounts of data they have been trained on [4]. Of these LLMs, Generative Pre-trained Transformer (GPT) LLMs, such as ChatGPT, have showcased unprecedented proficiency in the human language, coding, logic, reasoning, and other associated natural language tasks [33]. They excel at guiding complex conversations and are used in countless endeavours such as software engineering (SE) - as in AI for Software Engineering (AI4SE) [22] - and qualitative research [26].
Within the AI4SE field, the capabilities of LLMs are being explored in the software development, maintenance, and evolution [28, 30, 26, 22], namely LLM4SE. Accordingly, they find applications across various stages of the software development life cycle, including requirement analysis, software design, code implementation, testing, refactoring, defect detection, and maintenance [30]. Albeit the notable use of LLM in SE, to the best of our knowledge there are no studies that focus on the use of LLM in the ethical AI development. In qualitative research, LLMs are also gaining prominence, particularly as a supplement to tasks traditionally performed by humans, like analysing qualitative data [12, 32, 24]. Specifically, in the coding process - where qualitative data is organised and interpreted by assigning codes or labels to text segments or different data forms - LLMs have demonstrated significant utility [36, 32].
@@
Effective AI4SE should be trustworthy and synergistic with the practitioner’s workflow, otherwise “such AI4SE solutions risk becoming obstacles rather than facilitators" [22]. Currently, there is a growing need for context-dependent empirical studies that explore how trust in LLMs affects their adoption in software engineering tasks [22]. Furthermore, there remains a significant gap in research focused on the practical operationalisation of AI ethics, particularly through the application of LLMs. To date, no studies have attempted to explore the development of ethical AI-based systems using LLMs.
In this paper, we aim to explore trustworthiness in LLMs in the ethical AI development through an experiment. Our Research Question is: To what extent does the trustworthiness of LLM-based systems contribute to achieving ethically aligned AI-based systems? To this end, an empirical study was carried out, following the Design Science Research (DSR) method [13]. We identify techniques that can improve trustworthiness in LLMs, apply them to develop a prototype, use it in the context of the development of ethically aligned AI-based systems, and evaluate the output.
The evaluation is conducted through four different approaches. First, a thematic analysis is performed by an LLM, which identifies and categorizes key themes within the data. Second, we apply hierarchical clustering dendogram to group related text to compare and support the thematic analysis. Third, we perform an ablation study. Fourth, we run the source code produced to assess its functionality and correctness.
The contributions of this paper to the current state of knowledge on LLM4SE and ethics in AI are threefold. First, we showcase the effectiveness of various techniques in enhancing trustworthiness in LLM4SE. Second, we provide empirical insights into how these techniques can be leverage in developing ethically-aligned AI-based systems. Finally, by applying a multi-faceted evaluation approach - including thematic analysis, hierarchical clustering, and direct source code execution - we provide an initial framework for assessing the reliability and trustworthiness of LLMs in software engineering. With this experiment, we hope to shed light both on how to improve trustworthiness in LLM and on how to help practitioners develop ethical AI-based systems.
@@ 2.1 Large Language Models
LLMs have been pre-trained on vast amounts of text data, often scraped from the internet, allowing them to learn patterns and structures inherent in human language [14]. One of the key features of LLMs is their ability to generate contextually relevant and apparently coherent text across a wide range of topics [33]. These models can perform tasks such as language translation, text summarization, question answering, and creative writing [14]. They achieve this by leveraging the vast amounts of data they have been trained on to predict the next word or sequence of words in each context [10]. Some well-known examples of LLMs include GPT (Generative Pre-trained Transformer) models developed by OpenAI [27], particularly GPT-3.5 and GPT-4.
@@ 2.2 Trustworthiness in Large Language Models
Recent advancements and use of LLMs have raised concerns about their ethical implications [40, 19, 35, 20, 37]. Issues like information hallucination, biases, and the need for factual correctness are critical in determining the trustworthiness of these systems in real-world applications [18, 9]. Trustworthiness refers to the quality or attribute of being reliable, dependable, and deserving of trust. It encompasses several key components that establish trust in a person, organization, product, or system [37]. In the context of LLMs, trustworthiness largely refers to their reliability and ethical adherence to social norms [20, 37]. Notably, several studies have formulated taxonomies accompanied by evaluation methodologies, wherein the taxonomy acts as a measure for assessing trustworthiness, with the identified aspects serving as evaluation parameters. Among these efforts, the Holistic Evaluation of Language Models (HELM) devised by Liang et al. [19] introduces a comprehensive taxonomy consisting of seven metrics — accuracy, calibration, robustness, fairness, bias, toxicity, and efficiency — and applied this framework to analyze 30 different models. Notably, the authors stress the significance of adopting a multi-metric approach, enabling metrics prioritization beyond mere accuracy.
In a conspicuous study, partially funded by distinguished agencies such as NASA, Ebay, Amazon and Google, which also won the Outstanding Paper Award at NeurIPS’23, Wang et al. [40] provide eight trustworthiness perspectives of language models: toxicity, stereotype and bias, adversarial robustness, out-of-distribution robustness, privacy, robustness to adversarial demonstrations, machine ethics, lastly, fairness. The authors state that they only provided objective definitions for some of the perspectives coined - due to their intrinsic subjectivity, e.g., fairness toxicity - leaving the subjective exploration of model behaviors based on human understanding as future work. As scholars delve into characterizing trustworthiness within LLMs, a unified terminology for measuring it remains unclear. Furthermore, it is seen that there is no clear consensus on a framework to build aligned LLMs, nor clear assessment guidelines, both posing as unresolved issues [20].
2.3 Techniques to improve trustworthiness in LLM4SE
@@ 2.5 LLMs for Qualitative Analysis
Thus, LLMs can tackle the challenges of traditional qualitative research, such as the time-intensive nature of data analysis, limited generalizability, consistency issues, and personal subjective biases [1, 12]. Hamilton et al. [12] argue that both humans and LLMs have identified unique and overlapping themes from interview data, indicating the potential for recognizing patterns and themes in qualitative data.
@@ 3 Methodology
In this study, we employed an adapted DSR [13] approach to explore and enhance trustworthiness in LLMs for software engineering tasks in four phases: 1) exploration, 2) prototyping, 3) evaluation, and 4) communication of results. The fourth phase will be approached in Section 4.
@@ 3.1 Exploring
Initially, we explored existing literature and found the need and motivation to improve trustworthiness in AI4SE, presented in Section 2.2, and most notably in the work of David Lo [22]. Furthermore, we identified techniques that have been recently proposed in the literature to improve trustworthiness in AI systems, particularly in the context of software engineering, i.e., AI4SE. In sum, the identified techniques described in Section 2.3 provide an LLM-based multi-agent systems approach, characterised by: a) a combination of agents, b) distinct roles (specialized agents), c) structured conversation, d) multiple rounds of debate.
These techniques were then synthesised and integrated into the development of a prototype system designed to assist in software engineering tasks, in particular in the creation of ethically-aligned AI-based systems.
3.2 Prototype: LLM-BMAS
From our findings, we devised a prototype that is an LLM-Based multi-agent system (henceforth, called LLM-BMAS). The LLM-BMAS is a Python script that consists of three agents making use of GPT-4o-mini model through OpenAI API. They have different specialized roles and their conversation is structured in the same way. Specifically, there are two senior Python developers and one AI ethicist. As a team, they receive an input - a project description (PD)- that they should work on collaboratively to implement, through five rounds of discussion.
The conversational pattern assigned to them is as follows: Agent 1 should start the conversation; Agent 2 receives this as input to work on; similarly Agent 3 receives both conversations appended to work on. This defines a round of conversation. In other words, the conversation history is always appended, serving as input to the next agent until the number of rounds is reached. Figure 1 presents an overview of our prototype.
@@ 3.2 Prototype: LLM-BMAS
The agents engage in a structured conversation, spanning five rounds, to discuss and collaborate on the project. However, due to the nature of LLMs, each new run produces a different software for the same input [31, 22], i.e., using the same system configuration for data generation generates a different debate and source code for each new invoke, in addition to a different thematic analysis. Thus, in order to enrich our analysis, we ran the system three different times to obtain distinct output.
3.3 Evaluation
To evaluate the effectiveness of the prototype, we applied a multi-faceted evaluation strategy. First, we conducted a thematic analysis using an LLM to automatically identify and categorize key themes within the generated output. This provided insights into the model’s alignment with ethical and functional requirements. Second, we employed hierarchical clustering to organize and group-related text segments, supporting and validating the findings of the thematic analysis by uncovering underlying patterns and relationships in the data. Then, both thematic analysis results and hierarchical clustering for each project description is compared and analysed. Finally, an ablation study is conducted following similar steps. Figure 2 presents a visualisation for the evaluation methodology devised.
An ablation study is useful to validate research results and consists in assessing performance of an AI system by removing components [23, 25]. In this case, we will remove the LLM-BMAS prototype and use as a baseline only the interaction with OpenAI ChatGPT user interface.
@@ 3.3 Evaluation
As discussed, each run of the prototype produces a different result for the same input, thus, we run the system three different times to obtain distinct output (yielding the raw output, e.g., $O1_{1}$, $O1_{2}$, $O1_{3}$ for the first project description).
Then, the collected data is analysed through the use of ChatGPT GPT-4o with a specific custom instruction acting as qualitative analyst. Each of the raw output produces a partial thematic analysis (e.g., $B2_{1}$, $B2_{2}$, $B2_{3}$ for the second project description). Consequently, the partial thematic analysis for each project description is summarised into one thematic analysis through the use of the same GPT-4o, yielding $TA$, $TB$ and $TC$.
In order to provide a hierarchical clustering, iRaMuTeQ software was used, which runs on top of R. The input for the generation of each hierarchical clustering is the append of the raw output of each project description (e.g., $O3_{1}$, $O3_{2}$, $O3_{3}$ together generates the hierarchical clustering for the third project description.
With the aim of understanding the benefits of this LLM-BMAS approach, an ablation study is also performed, in which a project description is prompted to a regular ChatGPT-4o using OpenAI user interface, without any of the techniques found. For example, for project description 1, we have only one output ($AS_{1}$), that servers as input for creation of its thematic analysis also with ChatGPT-4o, and for the iRaMuTeQ analysis, yielding $TA_{A}S$ and $HCD_{A}S$ respectively.
Finally, we performed a practical evaluation of the prototype by running the source codes generated by the LLM. This step is for evaluating the system’s practical utility in real-world software development scenarios. Our methodology offers a initial framework for assessing the trustworthiness and reliability of LLMs in the software engineering domain.
The results and discussions arising from the comparison between the different thematic analysis and hierarchical clustering, coupled with the ablation study and source code execution is presented next.

### v2

@@ Abstract
AI-based systems, including Large Language Models (LLM), impact millions by supporting diverse tasks but face issues like misinformation, bias, and misuse. AI ethics is crucial as new technologies and concerns emerge, but objective, practical guidance remains debated. This study examines the use of LLMs for AI ethics in practice, assessing how LLM trustworthiness-enhancing techniques affect software development in this context. Using the Design Science Research (DSR) method, we identify techniques for LLM trustworthiness: multi-agents, distinct roles, structured communication, and multiple rounds of debate. We design a multi-agent prototype LLM-MAS, where agents engage in structured discussions on real-world AI ethics issues from the AI Incident Database. We evaluate the prototype across three case scenarios using thematic analysis, hierarchical clustering, comparative (baseline) studies, and running source code. The system generates approximately 2,000 lines of code per case, compared to only 80 lines in baseline trials. Discussions reveal terms like bias detection, transparency, accountability, user consent, GDPR compliance, fairness evaluation, and EU AI Act compliance, showing this prototype ability to generate extensive source code and documentation addressing often overlooked AI ethics issues. However, practical challenges in source code integration and dependency management may limit its use by practitioners.
@@ 1 Introduction
Artificial Intelligence (AI) is emerging as a transformative force, reshaping industries, economies and everyday life. Despite its rapid development and adoption, a number of negative reports on its use highlight the importance of adhering to ethical norms and principles [1]. Several ethical guidelines are available with plenty of ethical principles providing high level and abstract guidance for developers and stakeholders on AI ethics [2]. These are important, but there is a lack of practical guidance for developers to operationalise ethics in AI [2]. As new AI-based technologies emerge, ethics in AI will also become increasingly critical, such as the latest breakthrough: Large Language Models (LLMs).
LLMs are becoming ubiquitous and have significant impact on decision-making processes and human interactions [3, 4]. LLMs are advanced AI algorithms capable of generating, interpreting, and predicting text based on vast amounts of data they have been trained on [5]. Of these LLMs, Generative Pre-trained Transformer (GPT) LLMs, such as GPT-4o, have showcased unprecedented proficiency in the human language, coding, logic, reasoning, and other associated natural language tasks [4]. They excel at guiding complex conversations and are used in countless endeavours such as software engineering (SE) - as in AI for Software Engineering (AI4SE) [6] - and qualitative research [7].
Within the AI4SE field, the capabilities of LLMs are being explored in the software development, maintenance, and evolution [8, 9, 7, 6], namely LLM4SE. Accordingly, they find applications across various stages of the software development life cycle, including requirement analysis, software design, code implementation, testing, refactoring, defect detection, and maintenance [9]. Albeit the notable use of LLM in SE, to the best of our knowledge there are no studies that focus on the use of LLM in the ethical AI development. In qualitative research, LLMs are also gaining prominence, particularly as a supplement to tasks traditionally performed by humans, e.g., analysing qualitative data [10, 11, 12]. Specifically, in the coding process - where qualitative data is organised and interpreted by assigning codes or labels to text segments or different data forms - LLMs have demonstrated significant utility [13, 11].
@@
Effective AI4SE should be trustworthy and synergistic with the practitioner’s workflow, otherwise “such AI4SE solutions risk becoming obstacles rather than facilitators" [6]. Currently, there is a growing need for context-dependent empirical studies that explore how trust in LLMs affects their adoption in software engineering tasks [6]. Furthermore, there remains a significant gap in research focused on the practical operationalisation of AI ethics, particularly through the application of LLMs. To date, no studies have attempted to explore the development of ethical AI-based systems using LLMs. We aim to explore trustworthiness-enhancing techniques in LLMs in the ethical AI development context.
The following Research Question guides this study:
  • RQ: To what extent can trustworthiness-enhancing techniques in LLM-based systems support the development of ethically aligned AI software?
To this end, we employed an adapted single cycle of the Design Science Research (DSR) method [17] to identify trustworthiness-enhancing techniques, build a prototype, evaluate it through three case studies and communicate the results. We identified four trustworthiness-enhancing techniques for LLM-based systems in the literature and implemented them: (1) multi-agent collaboration, (2) specialized roles, (3) structured communication, and (4) multiple rounds of debate. Furthermore, we evaluated the prototype through three case studies based on real-world AI incidents in recruitment, deepfake detection, and image classification. Each case is analyzed using (1) LLM for thematic analysis, (2) iRaMuTeQ for hierarchical clustering, (3) a comparative baseline with standard ChatGPT GPT-4o prompts, and (4) source code execution to verify whether the generated code runs, without assessing its runtime behaviour or correctness in depth. The prototype can generate around 2,000 lines of code (with proper documentation), encompassing themes such as bias detection, transparency, accountability, user consent and GDPR compliance, fairness evaluation and compliance with the EU AI Act. The hierarchical clustering analysis revealed the ability of the prototype to generate outputs related to legal aspects of ethics in AI through words such as: bias_type, evaluate_bias, bias_score. The baseline study produced around 80 lines of output without source code while covering considerably less AI ethical issues. Therefore, the evaluations conducted suggest that the trustworthiness-enhancing techniques employed can be beneficial in the AI ethical development context. However, practical challenges in source code integration and dependency management may limit its use by practitioners, e.g., deprecated dependencies and modularity. We hope to shed light both on how to improve trustworthiness in LLM and on how to help practitioners to operationalise ethical principles in the development of AI-based systems.
@@ 2.1 Large Language Models
LLMs have been pre-trained on vast amounts of text data, often scraped from the internet, allowing them to learn patterns and structures inherent in human language [18]. One of the key features of LLMs is their ability to generate contextually relevant and apparently coherent text across a wide range of topics [4]. These models can perform tasks such as language translation, text summarization, question answering, and creative writing [18]. They achieve this by leveraging the vast amounts of data they have been trained on to predict the next word or sequence of words in each context [19]. Some well-known examples of LLMs include Generative Pre-trained Transformer (GPT) models developed by OpenAI [20], particularly GPT-4o and o3.
@@ 2.2 Trustworthiness in Large Language Models
Recent advancements and use of LLMs have raised concerns about their ethical implications [21, 22]. Issues like information hallucination, biases, and the need for factual correctness are critical in determining the trustworthiness of these systems in real-world applications [23, 24]. Trustworthiness refers to the quality or attribute of being reliable, dependable, and deserving of trust. It encompasses several key components that establish trust in a person, organization, product, or system [25]. In the context of LLMs, trustworthiness largely refers to their reliability and ethical adherence to social norms [3, 25]. Notably, several studies have formulated taxonomies accompanied by evaluation methodologies, wherein the taxonomy acts as a measure for assessing trustworthiness, with the identified aspects serving as evaluation parameters. Among these efforts, the Holistic Evaluation of Language Models (HELM) devised by Liang et al. [22] introduces a comprehensive taxonomy consisting of seven metrics — accuracy, calibration, robustness, fairness, bias, toxicity, and efficiency — and applied this framework to analyze 30 different models. Notably, the authors stress the significance of adopting a multi-metric approach, enabling metrics prioritization beyond mere accuracy.
Wang et al. [21] provide eight trustworthiness perspectives of language models: toxicity, stereotype and bias, adversarial robustness, out-of-distribution robustness, privacy, robustness to adversarial demonstrations, machine ethics, lastly, fairness. The authors state that they only provided objective definitions for some of the perspectives coined - due to their intrinsic subjectivity, e.g., fairness toxicity - leaving the subjective exploration of model behaviors based on human understanding as future work. As scholars delve into characterizing trustworthiness within LLMs, a unified terminology for measuring it remains unclear. Furthermore, it is seen that there is no clear consensus on a framework to build aligned LLMs, nor clear assessment guidelines, both posing as unresolved issues [3].
2.3 Trustworthiness enhancing techniques in LLM4SE
@@ 2.5 LLMs for Qualitative Analysis
Thus, LLMs can tackle the challenges of traditional qualitative research, such as the time-intensive nature of data analysis, limited generalizability, consistency issues, and personal subjective biases [37, 10]. Hamilton et al. [10] argue that both humans and LLMs have identified unique and overlapping themes from interview data, indicating the potential for recognizing patterns and themes in qualitative data. Drápal et al. [36] proposed a framework for collaboration between legal experts and LLM for performing thematic analysis. The LLM used could identify themes that would probably be missed due to human error. Also, they discuss that LLM can perform the initial coding of the data with reasonable quality. De Paoli [39] states that “there is no denying that the model can infer codes and themes and even identify patterns which researchers did not consider relevant and contribute to better qualitative analysis.”
Differently from the aforementioned approaches, our study aims to provide case studies using LLMs automating the whole thematic analysis process, as well as implementing the trustworthiness-enhancing techniques found to propose a prototype that generates source code in the AI ethical development context. For the best of our knowledge, there are no works in the literature that aims to use LLM to operationalise AI ethical principles.
@@ 3 Methodology
In this study, we used an adapted DSR approach [17] to explore and enhance the trustworthiness of LLMs for software engineering tasks. This approach involved four phases: (1) exploration, (2) prototyping, (3) evaluation through case studies and (4) communication of results. Due to the limited scope of this initial case study, this was done in one single cycle. The evaluation phase consists of applying the LLM-MAS prototype to three case scenarios, each derived from real-world AI incidents, and analysing the generated outcomes using thematic analysis, hierarchical clustering, comparative baselines, and source code execution. These case studies serve to examine how the identified trustworthiness-enhancing techniques impact the development of ethically aligned AI-based systems in practice. This Section will cover phases 1 to 3, while phase 4 will be discussed in Section 4.
@@ 3.1 Exploring
Initially, we explored the existing literature and identified the need to improve trustworthiness in AI4SE, as presented in Section 2.2. This need is most clearly demonstrated in the work of David Lo [6]. Furthermore, we identified recently proposed techniques in the literature to enhance trustworthiness in LLMs, particularly in the context of software engineering (i.e. LLM4SE). In summary, the trustworthiness-enhancing techniques identified and presented in Section 2.3 provide the basis for our conceptual approach, which is visualised in Table 1.
Table 1: Trustworthiness-enhancing techniques adopted
Trustworthiness-enhancing technique Description
Multi-agent collaboration Multiple agents interact on a task
Specialized roles Agents assume roles, e.g., Developer or Ethicist
Structured communication Predefined message passing structure
Multiple debate rounds Iterative dialogue to refine results
Next, we will present a detailed overview of the technical implementation of the prototype.
3.2 Prototype: LLM-MAS Technical Implementation
We devised a prototype of an LLM-based Multi-Agent System (henceforth referred to as the LLM-MAS). The LLM-MAS is a Python script consisting of three agents that make use of the GPT-4o-mini model via the OpenAI API. The model was selected on the basis of its cost-efficiency, accessibility, and strong reasoning performance, which rendered it a practical choice for iterative experimentation involving multiple agents. Rather than being equipped with memory or tool access, the agents are LLMs with custom instructions in the form of system prompts that describe their roles and the communication structure. Consequently, they have different specialised roles, and their conversations are structured in the same way. Specifically, there are two senior Python developers and one AI ethicist. The conversation structure comprises: reply, reflection, code, and critique. The agent prompts are available in the Zenodo: agent1_prompt_PythonDev.txt, agent2_prompt_PythonDev.txt and agent3_prompt_AIEthicist.txt.
The workflow for the agents is sequential. A round is defined as a sequence of conversations between agents 1, 2 and 3. The conversation history is appended after each agent output, and each agent receives the entire conversation history as an input. The number of rounds is modifiable. We selected five rounds per project to allow agents enough interaction to develop and refine ideas collaboratively, while keeping the output concise and manageable for analysis. Figure 1 presents an overview of our prototype.
@@ 3.2 Prototype: LLM-MAS Technical Implementation
The agents engage in a structured conversation, through multiple rounds, to discuss and collaborate on a project. That is, as a team, they receive an input – a project description (PD) – that they should work on collaboratively to implement, through structured, iterative discussions.
3.3 Evaluation: Case Studies
We begin with an overview of the evaluation approach, followed by an in-depth explanation of each component.
To evaluate the effectiveness of the prototype, we conducted a case study-based evaluation across three distinct scenarios inspired by real-world AI incidents. Each case study involved applying the LLM-MAS prototype to a different project description, enabling us to assess how the system behaved in various ethical and technical contexts. For each case, we employed a multi-faceted evaluation strategy. First, we performed a thematic analysis using an LLM to automatically identify and categorise key ethical and technical themes in the output. This helped us to assess the system’s ability to address ethical requirements. Secondly, hierarchical clustering (via iRaMuTeQ) was used to group related text segments, providing an additional layer of validation by highlighting patterns in the generated content. Thirdly, we conducted a comparative baseline study, prompting the same project descriptions directly to GPT-4o in ChatGPT interface (without any multi-agent configuration), to evaluate the added value of the proposed trustworthiness-enhancing techniques. Figure 2 presents an overview of the evaluation methodology approached.
@@ 3.3 Evaluation: Case Studies
Due to the probabilistic nature of LLMs, running the same input multiple times can yield different outputs. This means that even when using the same system configuration, each run may produce a different conversation, source code, and thematic analysis. Prior work has similarly observed this variability in multi-agent LLM systems [27, 6]. Thus, in order to enrich our analysis, we ran the system three different times to obtain distinct output (yielding the raw output, e.g., $O1_{1}$, $O1_{2}$, $O1_{3}$ for the first project description).
Then, the collected data is analysed through the use of ChatGPT GPT-4o with a specific custom instruction acting as qualitative analyst, available in the file Custom_instructions_Qualitative_GPT.txt in Zenodo. The prompt used as a header for the thematic analysis in available in file Prompt_Perform_thematic_analysis.txt. Each of the raw output produces a partial thematic analysis (e.g., $B2_{1}$, $B2_{2}$, $B2_{3}$ for the second project description). Consequently, the partial thematic analysis for each project description is summarised into one thematic analysis through the use of the same GPT-4o, yielding $TA$, $TB$ and $TC$. The prompt serving as header for the merging of the thematic analysis is available in file Prompt_Merging_Thematic_Analysis.txt, equipped with the same custom instruction.
We used Descending Hierarchical Classification (DHC) from iRaMuTeQ v0.7-alpha running in R v4.2. The input for the generation of each hierarchical clustering is the append of the raw output of each project description (e.g., $O3_{1}$, $O3_{2}$, $O3_{3}$ together generates the hierarchical clustering for the third project description.
With the aim of understanding the benefits of this LLM-MAS approach, a baseline study is also performed, in which a project description is prompted to a regular ChatGPT-4o using OpenAI user interface, without any of the techniques found. For example, for project description 1, we have only one output ($AS_{1}$), that servers as input for creation of its thematic analysis also with ChatGPT-4o, and for the iRaMuTeQ analysis, yielding $TA_{A}S$ and $HCD_{A}S$ respectively.
Finally, we performed a practical evaluation of the prototype by running the source codes generated by the LLM. This step is for evaluating the system’s practical utility in real-world software development scenarios. Our methodology offers a initial framework for assessing the trustworthiness and reliability of LLMs in the software engineering domain. The results and discussions arising from the comparison between the different thematic analysis and hierarchical clustering, coupled with the baseline study and source code execution is presented next.